# Gaussian Process Tilted Nonparametric Density Estimation using Fisher Divergence Score Matching

### v1

@@ Abstract
We present three Fisher divergence (FD) minimization algorithms for learning Gaussian process (GP) based score models for lower dimensional density estimation problems. The density is formed by multiplying a base multivariate normal distribution with an exponentiated GP refinement, and so we refer to it as a GP-tilted nonparametric density. By representing the GP part of the score as a linear function using the random Fourier feature (RFF) approximation, we show that all learning problems can be solved in closed form. This includes the basic and noise conditional versions of the Fisher divergence, as well as a novel alternative to noise conditional FD models based on variational inference (VI). Here, we propose using an ELBO-like optimization of the approximate posterior with which we derive a Fisher variational predictive distribution. The RFF representation of the GP, which is functionally equivalent to a single layer neural network score model with cosine activation, provides a unique linear form for which all expectations are in closed form. The Gaussian base also helps with tractability of the VI approximation. We demonstrate our three learning algorithms, as well as a MAP baseline algorithm, on several low dimensional density estimation problems. The closed-form nature of the learning problem removes the reliance on iterative algorithms, making this technique particularly well-suited to large data sets.
@@ I Introduction
Density estimation is a fundamental component of generative modeling. With deep neural networks, density estimation has been significantly improved using autoregressive models [1, 2, 3, 4], VAEs [5, 6], energy-based models [7], normalizing flows [8, 9, 10, 11] and GANs [12].
Parallel to these developments, score-based methods have emerged as a powerful learning approach for generative models [13]. Instead of optimizing the log-likelihood of the data, it uses the Fisher divergence between the model and data distributions to match the log gradients of the two distributions,
@@
[…]
The advantages of using Fisher divergence for score matching have been demonstrated for breaking the curse of dimensionality [18] and solving general inverse problems [19], such as compressed sensing MRI [20]. An additional benefit of using score matching is that the score function $s_{\theta}(x)$ does not need to correspond to a textbook density and so can be any neural network function, in effect making the approximation nonparmetric.
While clearly very powerful, these approaches also suffer from common limitations of learning complex networks of many parameters. In this paper, we focus on lower density estimation problems for which a simpler network is sufficient. In particular, we base our score function on a Gaussian process tilted density. Gaussian processes and tilted distributions have been shown to work well for nonparametric density estimation problems [21, 22, 23]. Kernel-based exponential family ideas have also been introduced for score matching, leading to iterative learning algorithms and deep representations [24, 25, 26, 27]. In contrast with these methods, our algorithms will have closed form solutions.
We propose a GP-tilted density estimator using the random Fourier feature approximation and learned it via Fisher divergence. Our selected GP is mathematically equivalent to a single layer fully connected neural network with randomized parameters; the only parameter to be learned is a linear layer vector with dimension equal to the approximation level of the GP. We derive a noise-conditional method and a novel variational predictive distribution for learning, both of which have closed form solutions as a result of our chosen score function. We evaluate our Fisher divergence-based algorithms, as well as MAP and kernel density estimation, on several low dimensional problems.
Our main contributions are two-fold:
@@
    Our model is equivalent to an exponentiated single layer neural network for which one parameter vector needs to be learned (on the order of 1000s of parameters). We present three different algorithms for this based on modifications of the Fisher divergence, all of which give closed form solutions for this vector, making the technique particularly well-suited for large data sets.
@@
    We propose a new approach to the Fisher divergence based on a variational approximation to the model posterior using an ELBO-like function and variational tempering. The corresponding predictive distribution on the data integrates out uncertainty in the model parameter rather than by adding noise to the data, giving an alternative to noise-conditional Fisher divergence methods for our model.
The paper is organized as follows: In Section 2, we provide a brief background on the Gaussian process and the Fisher divergence score matching methods that is relevant for our derivations. In Section 3, we present the GP-tilted density model and derive three different, closed-form solutions for learning its parameter vector. We then demonstrate the model on several low dimensional data sets in Section 4.
@@ II Background
We first briefly review background on Gaussian processes and Fisher divergence minimization that we use for our GP-tilted density estimation model and algorithms.
@@ II-A Gaussian Processes
Let $k(x,x^{\prime})$ be a kernel function between two vectors $x,x^{\prime}\in\mathbb{R}^{D}$. Then the function $f:\mathbb{R}^{D}\rightarrow\mathbb{R}$ is a zero-mean Gaussian process if, for any $N$ vectors $x_{1},\dots,x_{N}$, the corresponding function evaluations $[f(x_{1}),\dots,f(x_{N})]^{\top}\sim\mathcal{N}(0,K_{N})$, where $K_{N}(i,j)=k(x_{i},x_{j})$. We write $f(x)\sim\mathcal{GP}(0,k(x,x^{\prime}))$.
There are a large variety of kernel functions $k$ that can be used with the GP, and extensions can be made to other spaces and non-zero mean functions. The most popular kernel in general is the Gaussian kernel of the form $k(x,x^{\prime})=\lambda\exp(-\frac{1}{2\gamma^{2}}\|x-x^{\prime}\|^{2})$, which measures correlations based on proximity in the input space. For the Gaussian kernel, the following random Fourier feature (RFF) definition provides an asymptotically exact way of representing the underlying linear model of the GP in its reproducing kernel Hilbert space [28]. The equivalent structure of this representation to a single layer of a neural network has recently motivated other applications of this idea to traditional machine learning problems [29, 30].
@@
Let $S$ to be a positive integer, $\gamma>0$, and $x\in\mathbb{R}^{D}$. For $s=1,\dots,S$, generate $z_{s}\sim\mathcal{N}(0,I)$ and $c_{s}\sim\mathrm{Unif}(0,2\pi)$. The random Fourier feature mapping of the vector $x$ is defined to be
$\phi(x)=\sqrt{\frac{2}{S}}\left[\cos{\left(\frac{z_{1}^{\top}x}{\gamma}+c_{1}\right)},\dots,\cos{\left(\frac{z_{S}^{\top}x}{\gamma}+c_{S}\right)}\right]^{\top}$
For any $x,x^{\prime}\in\mathbb{R}^{D}$, $\phi(x)^{\top}\phi(x^{\prime})\rightarrow\exp(-\frac{1}{2\gamma^{2}}\|x-x^{\prime}\|^{2})$ as $S\rightarrow\infty$. Thus $\phi$ approximates the RKHS of the GP.
It follows that the marginal of $f_{\theta}(x)=\theta^{\top}\phi(x)$ over $\theta$, for $\theta\sim\mathcal{N}(0,\lambda^{-1}I)$, has distribution $\mathcal{GP}(0,k(x,x^{\prime}))$ with $k(x,x^{\prime})=\lambda^{-1}\phi(x)^{\top}\phi(x^{\prime})\rightarrow\lambda^{-1}\exp(-\frac{1}{2\gamma^{2}}\|x-x^{\prime}\|^{2})$ in the limit as $S\rightarrow\infty$.
II-B Fisher divergence for learning
Let $\mathcal{P}$ be the true unknown distribution of data from which we have i.i.d. samples, $x\sim\mathcal{P}$. Our goal is to approximate the density $p(x)$ corresponding to $\mathcal{P}$ with a parametric density $q_{\theta}(x)$. This is achieved by tuning parameters $\theta$ such that the two densities agree according to some measure of similarity. Maximum likelihood is one standard approach based on the KL-divergence,
@@ II-B Fisher divergence for learning
$\theta_{\textrm{FD}}=\arg\min_{\theta}\frac{1}{2}\mathbb{E}_{p}\|\nabla_{x}\ln p(x)-\nabla_{x}\ln q_{\theta}(x)\|^{2}.$ (2)
Where a score function, typically a neural network and written in this paper as $\nabla_{x}\ln q_{\theta}(x)$, is designed to match the expected gradients of the log data distribution. Through a derivation involving integration by parts, this is equivalent to an optimization problem of the form
$\quad\theta_{\textrm{FD}}$$\approx$ $\textstyle\arg\min_{\theta}~{}\frac{1}{2}\sum_{i=1}^{N}\big{\|}\nabla_{x}\ln q_{\theta}(x_{i})\big{\|}^{2}\qquad\qquad$
$\qquad\quad~{}~{}\textstyle+\sum_{i=1}^{N}\mathrm{trace}\big{(}\nabla_{xx}^{2}\ln q_{\theta}(x_{i})\big{)}.$
A Monte Carlo approximated is again made using samples $x\sim\mathcal{P}$. The Fisher score focuses on creating agreement between the derivatives of $q_{\theta}$ and the true distribution $p$ in high density regions. Equation II-B shows that this is equivalent to placing a sharp mode around each $x_{i}$ in $q_{\theta}$.
For the GP-tilted density estimator discussed in Section III, this divergence has the advantage of producing a closed form solution for $\theta$, in contrast to maximum likelihood (see Section III-A). As with deep generative models, the same practical issue of accurately learning the low density regions arises, which we address in two ways: In Section III-B we adopt the frequently used noise conditional representation that still has a closed-form solution, and in Section III-C we propose a variational approximation to the predictive distribution that also has a closed-form solution.
@@ III GP-Tilted Density Estimation
[…]
As follows from Gaussian process theory and basic analysis, $q(x)$ is a continuous density [31]. The multivariate Gaussian base measure is used to ensure integrability of the numerator in $\mathbb{R}^{D}$. The mean $\mu$ and covariance $\Sigma$ broadly define the region of the data, while the exponentiated GP provides a nonparametric refinement. Using the RFF representation of $f(x)$ in Definition II.2, this density can be approximated by
@@
where $\phi$ is the $S$-dimensional random mapping. MAP inference for this model is straightforward, which we present in Algorithm 1 for reference. Functionally, the score $\ln q_{\theta}(x)$ has the form of a single layer neural network using RFF parameters. We next show that this approximation also allows for easy inference using Fisher divergence minimization.
@@
6:$\begin{array}[]{rcl}\textstyle\psi_{2}&=&\sum_{j}\phi(\zeta_{j})w_{t}(j),\\ \textstyle\qquad w_{t}(j)&=&\exp\{\theta_{t}^{\top}\phi(\zeta_{j})\}\big{/}\sum_{j^{\prime}}\exp\{\theta_{t}^{\top}\phi(\zeta_{j^{\prime}})\}\end{array}$
@@
III-A Fisher Divergence Score Matching
$\mathcal{L}=\frac{1}{2}\sum_{i}\big{\|}\nabla_{x}\ln q_{\theta}(x_{i})\big{\|}^{2}+\sum_{i}\mathrm{tr}\big{(}\nabla_{xx}^{2}\ln q_{\theta}(x_{i})\big{)}.$ (6)
(We account for the prior on $\theta$ later.) The gradient and Hessian of $\ln q_{\theta}(x)$ in our case are
$\nabla_{x}\ln q_{\theta}(x)$ $=$ $\textstyle\sum_{s}\frac{1}{\gamma}\theta_{s}\phi^{\prime}_{s}(x)z_{s}-\Sigma^{-1}(x-\mu),$ (7)
$\nabla_{xx}^{2}\ln q_{\theta}(x)$ $=$ $\textstyle\sum_{s}\frac{1}{\gamma^{2}}\theta_{s}\phi^{\prime\prime}_{s}(x)z_{s}z_{s}^{\top}-\Sigma^{-1},$ (8)
where for $s=1,\dots,S$ we have defined
$\phi^{\prime}_{s}(x)$ $=$ $-\sqrt{2/S}\sin(z_{s}^{\top}x/\gamma+c_{s}),$
$\phi^{\prime\prime}_{s}(x)$ $=$ $-\sqrt{2/S}\cos(z_{s}^{\top}x/\gamma+c_{s}).$ (9)
As is evident, after plugging Eqs. 7 & 8 into Equation 6, the Fisher divergence is a quadratic-plus-linear function of $\theta$. Define the matrix $Z=[z_{1},\dots,z_{S}]^{\top}$ and for any square matrix $M$, let $\|ZM\|^{2}\equiv(\|Mz_{1}\|^{2},\dots,\|Mz_{S}\|^{2})^{\top}$ — in this subsection, $M=I$. Optimizing for $\theta$ requires the following gradient,
$\nabla_{\theta}\mathcal{L}$ $=$ $\textstyle\Big{(}\sum_{i=1}^{N}\frac{1}{\gamma^{2}}ZZ^{\top}\odot\phi^{\prime}(x_{i})\phi^{\prime}(x_{i})^{\top}\Big{)}\theta$ (10)
$\textstyle-\sum_{i=1}^{N}\frac{1}{\gamma}\phi^{\prime}(x_{i})\odot Z\Sigma^{-1}(x_{i}-\mu)$
$\textstyle+\sum_{i=1}^{N}\frac{1}{\gamma^{2}}\phi^{\prime\prime}(x_{i})\odot\|Z\|^{2},$
where $\odot$ indicates elementwise multiplication.
@@ III-A Fisher Divergence Score Matching
$\theta_{\mathrm{FD}}$ $=$ $\textstyle\left(\lambda\gamma^{2}I+ZZ^{\top}\odot\sum_{i}\phi^{\prime}(x_{i})\phi^{\prime}(x_{i})^{\top}\right)^{-1}\times$
$\textstyle\left(\sum_{i}\gamma\phi^{\prime}(x_{i})\odot Z\Sigma^{-1}(x_{i}-\mu)+\phi(x_{i})\odot\|Z\|^{2}\right).$
This involves inverting an $S\times S$ matrix, where $S$ is the number of random Fourier features used to approximate the Gaussian process. When large values of $S$ or other numerical issues present a computational challenge, conjugate gradients can be used to quickly solve for $\theta$ instead. We summarize parameter estimation for the basic GP-tilted density using Fisher divergence in Algorithm 2.
@@
1:Data $x\in\mathbb{R}^{D}$, kernel width $\gamma>0$, param $\lambda>0$, $S\in\mathbb{Z}_{+}$, base parameterss $\mu\in\mathbb{R}^{D}$, $\Sigma\in\mathbb{S}_{++}^{d}$
@@
$\begin{array}[]{ll}\Phi^{\prime}&=~{}\textstyle\sum_{i}\phi^{\prime}(x_{i})\phi^{\prime}(x_{i})^{\top}\\ \psi^{\prime}&=~{}\textstyle\sum_{i}\phi^{\prime}(x_{i})\odot Z\Sigma^{-1}(x_{i}-\mu)\\ \psi&=~{}\textstyle\sum_{i}\phi(x_{i})\end{array}$
@@
III-A1 Learning base measure parameters $\mu$ and $\Sigma$
The base Gaussian measure of $q_{\theta}(x)$ in Equation 5 is important because it defines the region of focus for the Gaussian process. When set too small, the base resists the GP’s ability to increase the density near data outside this region. When it is too large, it requires the GP to be learned well in regions of no data, which isn’t likely since the GP $\theta^{\top}\phi(x)$ reverts to something like evaluations of the random prior. This can randomly produce large positive values away from the data that dominate the overall density $q_{\theta}(x)$.
We propose two modifications to address this issue in Section III-B and Section III-C. Here we note that, when optimizing the basic objective function in Equation 6, we can also include optimization over $\mu$ and $\Sigma$ as follows: For $\mu=\arg\min_{\mu}\mathcal{L}$, the optimal value given $\theta$ and $\Sigma$ is

### v2

@@ Abstract
We propose a nonparametric density estimator based on the Gaussian process (GP) and derive three novel closed form learning algorithms based on Fisher divergence (FD) score matching. The density estimator is formed by multiplying a base multivariate normal distribution with an exponentiated GP refinement, and so we refer to it as a GP-tilted nonparametric density. By representing the GP part of the score as a linear function using the random Fourier feature (RFF) approximation, we show that optimization can be solved in closed form for the three FD-based objectives considered. This includes the basic and noise conditional versions of the Fisher divergence, as well as an alternative to noise conditional FD models based on variational inference (VI) that we propose in this paper. For this novel learning approach, we propose an ELBO-like optimization to approximate the posterior distribution, with which we then derive a Fisher variational predictive distribution. The RFF representation of the GP, which is functionally equivalent to a single layer neural network score model with cosine activation, provides a useful linear representation of the GP for which all expectations can be solved. The Gaussian base distribution also helps with tractability of the VI approximation and ensures that our proposed density is well-defined. We demonstrate our three learning algorithms, as well as a MAP baseline algorithm, on several low dimensional density estimation problems. The closed form nature of the learning problem removes the reliance on iterative learning algorithms, making this technique particularly well-suited to big data sets, since only sufficient statistics collected from a single pass through the data is needed.
@@ I Introduction
Density estimation is a fundamental aspect of generative modeling. With neural networks, density estimation has been significantly improved for generative models using autoregressive models [larochelle2011neural, germain2015made, van2016conditional, van2016wavenet], VAEs [kingma2013auto, rezende2014stochastic, zhang2025explainable], energy-based models [song2021train], normalizing flows [rezende2015variational, dinh2016density, kingma2016improved, kingma2018glow, chen2025entropy] and GANs [goodfellow2014generative]. Parallel to these developments, score-based methods have emerged as a powerful learning approach for generative models [song2019generative]. In place of the log-likelihood function, it uses the Fisher divergence between the model and data distributions to match the log gradients of the two distributions,
@@
The gradient of the log density approximation, $s_{\theta}(x)$, also called the score function [liu2016kernelized], matches those of the true data distribution. Often this $s_{\theta}(x)$ is a complex neural network with parameters $\theta$. Many techniques have been proposed to optimize this divergence [hyvarinen2005estimation, vincent2011connection, song2020sliced] and samples are generated from $p_{\theta}(x)\propto\exp\{\int s_{\theta}(x)dx\}$ with Langevin dynamics. To avoid an indefinite integral we will opt to write $\nabla_{x}s_{\theta}(x)$ in (1).
The advantages of using Fisher divergence for score matching have been demonstrated for breaking the curse of dimensionality [cole2024score] and solving general inverse problems [song2021solving], such as compressed sensing MRI [jalal2021robust]. An additional benefit of using score matching is that the score function $s_{\theta}(x)$ does not need to correspond to a textbook density and so any neural network model can be used, in effect making the density estimation nonparmetric. While therefore clearly very powerful, the common limitations of choosing and learning complex neural networks of many parameters arises.
In this paper, we focus on lower density estimation problems for which a simpler network is sufficient. In particular, we base our score function on a proposed Gaussian process tilted density. Gaussian processes (GP) and tilted distributions have both been shown to work well for nonparametric density estimation problems [hensman2014tilted, dutordoir2018gaussian, floto2023tilted]. Kernel-based exponential family ideas have also been introduced to the score matching framework, leading to iterative learning algorithms and deep representations [wenliang2019learning, wang2020wasserstein, pacchiardi2022score, tsymboi2022denoising]. In contrast with these methods, our density will lead to non-iterative algorithms that calculate closed form solutions.
To achieve this, we propose a GP-tilted density estimator using the random Fourier feature approximation and learn it via Fisher divergence minimization. Our selected GP is mathematically equivalent to a single layer fully connected neural network with randomized parameters; the only parameter to be learned is a linear layer vector with dimension equal to the approximation level of the GP. We derive a noise conditional method and a novel variational predictive distribution for learning, both of which have closed form solutions as a result of our chosen score function. We evaluate our Fisher divergence-based algorithms, as well as MAP and kernel density estimation, on several low dimensional problems.
The main features of our Tilted Gaussian Process (TGP) density approximation method are two-fold:
@@
    Our score function is equivalent to a single hidden layer neural network for which one parameter vector needs to be learned due to randomization of the hidden layer parameters. We present three different algorithms for this based on modifications of the Fisher divergence, all of which give closed form solutions for this vector, making the technique particularly well-suited for efficiently learning with big data through parallel computing.
@@
    We propose a new approach to Fisher divergence minimization based on an approximation to the model posterior using an ELBO-like function and variational tempering. A Fisher variational predictive distribution on the data integrates out uncertainty in the model parameter rather than by adding noise to the data, giving an alternative to noise conditional Fisher divergence methods for our model.
The paper is organized as follows: In Section 2, we provide a brief review of aspects of the Gaussian process and the Fisher divergence score matching methods that are relevant for our derivations. In Section 3, we present the GP-tilted density model (TGP) and derive three different, closed form solutions for learning its parameter vector. We then demonstrate the model on several low dimensional data sets in Section 4.
@@ II Background
In this section, we briefly review background material on Gaussian processes and Fisher divergence minimization useful for our GP-tilted density estimation model and algorithms.
@@ II-A Gaussian Processes
The Gaussian process (GP) is a core nonparametric function approximation technique in probabilistic modeling and machine learning [williams2006gaussian]. Its formal definition is as follows:
@@
Let $k(x,x^{\prime})$ be a kernel function between two vectors $x,x^{\prime}\in\mathbb{R}^{d}$. Then the function $f:\mathbb{R}^{d}\rightarrow\mathbb{R}$ is a zero-mean Gaussian process if, for any $N$ vectors $x_{1},\dots,x_{N}$, the corresponding function evaluations $f(x_{1}),\dots,f(x_{N})$ are distributed as $\mathcal{N}(0,K_{N})$, where $K_{N}(i,j)=k(x_{i},x_{j})$. We write $f(x)\sim\mathcal{GP}(0,k(x,x^{\prime}))$.
There are a large variety of kernel functions $k$ that can be used with the GP, and extensions can be made to other spaces and non-zero mean functions. The most popular kernel for machine learning applications is the Gaussian kernel of the form $k(x,x^{\prime})=\lambda\exp(-\frac{1}{2\gamma^{2}}\|x-x^{\prime}\|^{2})$, which measures correlations based on proximity in the input space. For the Gaussian kernel, the following random Fourier feature (RFF) definition provides an asymptotically exact way of representing the underlying linear model of the GP in its reproducing kernel Hilbert space [rahimi]. The equivalent structure of this representation to a single layer of a neural network has recently motivated other applications of this idea to traditional machine learning problems [paisley2022bayesian, zhang2024a, zhang2024b].
@@
Let $S$ to be a positive integer, $\gamma>0$, and $x\in\mathbb{R}^{d}$. For $s=1,\dots,S$, generate $z_{s}\sim\mathcal{N}(0,I)$ and $c_{s}\sim\mathrm{Unif}(0,2\pi)$. Define a random Fourier feature mapping $\phi$ of the vector $x$ as
$\phi(x)=\sqrt{\frac{2}{S}}\left[\cos{\left(\frac{z_{1}^{\top}x}{\gamma}+c_{1}\right)},\dots,\cos{\left(\frac{z_{S}^{\top}x}{\gamma}+c_{S}\right)}\right]^{\top}.$
Then, for any $x,x^{\prime}\in\mathbb{R}^{d}$, $\phi(x)^{\top}\phi(x^{\prime})\rightarrow\exp(-\frac{1}{2\gamma^{2}}\|x-x^{\prime}\|^{2})$ as $S\rightarrow\infty$. Thus $\phi$ approximates the RKHS of the GP.
It follows from the marginal representation of the Gaussian process [williams2006gaussian] that, if $\theta\sim\mathcal{N}(0,\lambda^{-1}I)$, then the random function $f_{\theta}(x)=\theta^{\top}\phi(x)$ is distributed as $\mathcal{GP}(0,k(x,x^{\prime}))$ with $k(x,x^{\prime})=\lambda^{-1}\phi(x)^{\top}\phi(x^{\prime})$. Since this dot product converges to $\exp(-\frac{1}{2\gamma^{2}}\|x-x^{\prime}\|^{2})$ as $S\rightarrow\infty$, the approximation holds.
II-B Fisher Divergence for Parameter Learning
Let $\mathcal{P}$ be the true unknown distribution of data from which we have i.i.d. samples, $x\sim\mathcal{P}$. Our goal is to approximate the density $p(x)$ corresponding to $\mathcal{P}$ with another density $q_{\theta}(x)$. This is achieved by tuning parameters $\theta$ such that the two densities agree according to some measure of similarity. Maximum likelihood is one standard approach based on the KL-divergence,
@@ II-B Fisher Divergence for Parameter Learning
$\theta_{\textrm{FD}}=\arg\min_{\theta}\tfrac{1}{2}\mathbb{E}_{p}\left[\|\nabla_{x}\ln p(x)-\nabla_{x}\ln q_{\theta}(x)\|^{2}\right],$ (2)
where a score function, typically a neural network and written here as $\nabla_{x}\ln q_{\theta}(x)$, is designed to match the gradients of the log data distribution with emphasis on more probable regions. Through a derivation involving integration by parts, this is equivalent to an optimization problem of the form
$\quad\theta_{\textrm{FD}}\hskip-5.0pt$ $\approx$ $\hskip-5.0pt\arg\min_{\theta}~\tfrac{1}{2}\sum\nolimits_{i=1}^{N}\big\|\nabla_{x}\ln q_{\theta}(x_{i})\big\|^{2}\qquad\qquad$
$\qquad\quad\,+\sum\nolimits_{i=1}^{N}\mathrm{trace}\big(\nabla_{xx}^{2}\ln q_{\theta}(x_{i})\big).$
A Monte Carlo approximated is again made using i.i.d. samples $x\sim\mathcal{P}$. As mentioned, the Fisher score focuses on creating agreement between the derivatives of $q_{\theta}$ and the true distribution $p$ in high density regions. Equation II-B shows that this is done with $q_{\theta}$ by placing a mode around each $x_{i}$.
For the GP-tilted density estimator discussed next, the Fisher divergence has the advantage of producing a closed form solution for $\theta$, in contrast to maximum likelihood (see Section III-A). As observed with deep generative models, the same practical issue of accurately learning the low density regions arises, which we address in two ways: In Section III-B we adopt the frequently used noise conditional representation, which still has a closed form solution for our model, and in Section III-C we propose a variational approximation to the predictive distribution that also has a closed form solution.
@@ III GP-Tilted Density Estimation
$q(x)=\frac{\exp\{{f(x)\}}\mathcal{N}(x|\mu,\Sigma)}{\int\exp\{{f(x)\}}\mathcal{N}(x|\mu,\Sigma)\,dx}.$ (4)
As follows from Gaussian process theory and basic analysis, $q(x)$ is a continuous density [williams2006gaussian]. We use a multivariate Gaussian base measure to ensure integrability of the numerator in $\mathbb{R}^{d}$. The mean $\mu$ and covariance $\Sigma$ broadly define the region of interest over the data, while the exponentiated GP provides a nonparametric refinement. Using the RFF representation of $f(x)$ in Definition II.2, this density can be approximated by
@@
where $\phi$ is the $S$-dimensional RFF mapping. MAP inference for this model is fairly straightforward, which we present in Algorithm 1 for reference. Functionally, the score $\ln q_{\theta}(x)$ has the form of a single layer neural network using RFF parameters. In Figure 1 we show a simplified flowchart of the TGP density estimator, connecting it with a single layer neural network structure. We next show that this approximation also allows for easy inference using Fisher divergence minimization.
Figure 1: A simplified flowchart of the GP-tilted density. The input to the exponential function is the Gaussian process approximated by random Fourier features, which turns the GP into a single layer neural network. The only learnable parameter is the vector $\theta\in\mathbb{R}^{S}$. (We set $S=1000$ in our experiments.) A broad, predefined Gaussian based distribution determines the region of interest for the density to ensure integrability. In Algorithms 2-4, we present novel closed form solutions for learning $\theta$ and constructing predictive distributions based on the Fisher divergence score matching objective function.
III-A Fisher Divergence Score Matching
$\mathcal{L}=\tfrac{1}{2}\sum\nolimits_{i}\big\|\nabla_{x}\ln q_{\theta}(x_{i})\big\|^{2}+\sum\nolimits_{i}\mathrm{tr}\big(\nabla_{xx}^{2}\ln q_{\theta}(x_{i})\big).$ (6)
(We account for the prior on $\theta$ later.) The gradient and Hessian of $\ln q_{\theta}(x)$ for the TGP density are
$\nabla_{x}\ln q_{\theta}(x)\hskip-5.0pt$ $=$ $\hskip-5.0pt\sum\nolimits_{s}\tfrac{1}{\gamma}\theta_{s}\phi^{\prime}_{s}(x)z_{s}-\Sigma^{-1}(x-\mu),$ (7)
$\nabla_{xx}^{2}\ln q_{\theta}(x)\hskip-5.0pt$ $=$ $\hskip-5.0pt\sum\nolimits_{s}\tfrac{1}{\gamma^{2}}\theta_{s}\phi^{\prime\prime}_{s}(x)z_{s}z_{s}^{\top}-\Sigma^{-1},$ (8)
where for $s=1,\dots,S$ we have defined
$\phi^{\prime}_{s}(x)\hskip-3.0pt$ $=$ $\hskip-3.0pt-\sqrt{2/S}\sin(z_{s}^{\top}x/\gamma+c_{s}),$
$\phi^{\prime\prime}_{s}(x)\hskip-3.0pt$ $=$ $\hskip-3.0pt-\sqrt{2/S}\cos(z_{s}^{\top}x/\gamma+c_{s}).$ (9)
After plugging Eqs. 7 & 8 into Equation 6, the Fisher divergence is observed to be a quadratic-plus-linear function of $\theta$, which has an analytic optimal solution. Define the matrix $Z=[z_{1},\dots,z_{S}]^{\top}$ and for any square symmetric matrix $M$, let $\|ZM\|^{2}\equiv(\|Mz_{1}\|^{2},\dots,\|Mz_{S}\|^{2})^{\top}$ — in this subsection, $M=I$. Optimizing $\theta$ involves working with the gradient
$\nabla_{\theta}\mathcal{L}$ $=$ $\Big(\sum\nolimits_{i=1}^{N}\tfrac{1}{\gamma^{2}}ZZ^{\top}\odot\phi^{\prime}(x_{i})\phi^{\prime}(x_{i})^{\top}\Big)\theta$ (10)
$-\sum\nolimits_{i=1}^{N}\tfrac{1}{\gamma}\phi^{\prime}(x_{i})\odot Z\Sigma^{-1}(x_{i}-\mu)$
$+\sum\nolimits_{i=1}^{N}\tfrac{1}{\gamma^{2}}\phi^{\prime\prime}(x_{i})\odot\|Z\|^{2},$
where $\odot$ indicates elementwise multiplication.
@@ III-A Fisher Divergence Score Matching
[…]
@@
$\theta_{\mathrm{FD}}\hskip-5.0pt$ $=$ $\hskip-5.0pt\left(\lambda\gamma^{2}I+ZZ^{\top}\odot\sum\nolimits_{i}\phi^{\prime}(x_{i})\phi^{\prime}(x_{i})^{\top}\right)^{-1}\times$
$\hskip-5.0pt\left(\sum\nolimits_{i}\gamma\phi^{\prime}(x_{i})\odot Z\Sigma^{-1}(x_{i}-\mu)+\phi(x_{i})\odot\|Z\|^{2}\right).$
This involves inverting an $S\times S$ matrix, where $S$ is the number of random Fourier features used to approximate the Gaussian process. When large values of $S$ or other numerical issues present a computational challenge, conjugate gradients can be used to quickly find $\theta_{\mathrm{FD}}$ instead. We summarize parameter estimation for the basic GP-tilted density using Fisher divergence in Algorithm 2.
@@
1:Data $x\in\mathbb{R}^{D}$, kernel width $\gamma>0$, param $\lambda>0$, $S\in\mathbb{Z}_{+}$, base parameters $\mu\in\mathbb{R}^{D}$, $\Sigma\in\mathbb{S}_{++}^{d}$
@@
$\begin{array}[]{ll}\Phi^{\prime}&\hskip-5.0pt=~\textstyle\sum_{i}\phi^{\prime}(x_{i})\phi^{\prime}(x_{i})^{\top}\\ \psi^{\prime}&\hskip-5.0pt=~\textstyle\sum_{i}\phi^{\prime}(x_{i})\odot Z\Sigma^{-1}(x_{i}-\mu)\\ \psi&\hskip-5.0pt=~\textstyle\sum_{i}\phi(x_{i})\end{array}$
@@
III-A1 Learning $\mu$ and $\Sigma$ of base measure
The base Gaussian measure of $q_{\theta}(x)$ in Equation 5 is important because it defines the region of focus for the Gaussian process. When set too small, the base resists the GP’s ability to increase the density near data outside this region. When it is too large, it requires the GP to learn to suppress regions of no data, which is challenging because the GP $\theta^{\top}\phi(x)$ reverts to something like an evaluation of the random prior. This can randomly produce large positive values away from the data that dominate the overall density $q_{\theta}(x)$. We discuss two modifications to the Fisher divergence to address this issue in Section III-B and Section III-C.
For the basic Fisher divergence, one option is to set $\mu$ and $\Sigma$ to the empirical mean and scaled covariance of the data, where the scaling is done using cross-validation. We also note an option of performing coordinate descent over $\theta$, according to Equation 6, as well as $\mu$ and $\Sigma$ as follows: For $\mu=\arg\min_{\mu}\mathcal{L}$, the optimal value given $\theta$ and $\Sigma$ is